OpenStudy (quickstudent):

I'm multiplying and dividing functions, could someone just check my work for me?

OpenStudy (quickstudent):

Thanks, I'll type them in... this is the first one. f(x) = 4x g(x) = x + 5 f(x) * g(x) = ? 4x(x + 5) = 4x * x + 4x * 5 = 4x^2 + 4x * 5 = 4x^2 + 20x

OpenStudy (anonymous):

Excellent.

OpenStudy (quickstudent):

Thanks

OpenStudy (quickstudent):

Could you check a few more for me please?

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

Sure

OpenStudy (quickstudent):

f(x) = 2x + 1 g(x) = 3x + 2 f(3)/g(2) = ? (2 * 3 + 1)/(3 * 2 + 2) = (6 + 1)/(3 * 2 + 2) = 7/(3 * 2 + 2) = 7/(6 + 2) = 7/8

OpenStudy (anonymous):

Correct. You're really good at this.

OpenStudy (quickstudent):

Thanks:) A few more now... f(x) = x^2 - 5 g(x) = 3x - 1 f(4)*g(-1) = ? (4^2 - 5) * (3 * (-1) - 1) = (16 - 5) * (3 * (-1) - 1) = 11 * (3 * (-1) - 1) = 11 * (-3 - 1) = 11 * (-4) = -44

OpenStudy (anonymous):

Right again

OpenStudy (quickstudent):

:D Here's the next one f(x) = x - 4 g(x) = 3x + 6 f(5) + g(1)/f(6) - g(-2) = ? {(5 - 4) + (3 * 1 + 6)}/{(6 - 4) - (3 * (-2) + 6)} = {1 + (3 * 1 + 6)}/{(6 - 4) - (3 * (-2) + 6)} = {1 + (3 + 6)}/{(6 - 4) - (3 * (-2) + 6)} = {1 + 9}/{(6 - 4) - (3 * (-2) + 6)} = 10/{(6 - 4) - (3 * (-2) + 6)} = 10/{2 - (3 * (-2) + 6)} = 10/{2 - (-6 + 6)} = 10/{2 - 0} = 10/2 = 5

OpenStudy (anonymous):

Excellent work

OpenStudy (quickstudent):

Thank you so much. Just 4 more now. f(x) = 3x g(x) = 3x - 6 f(x) * g(x) = ? 3x(3x - 6) = 3x * 3x - 3x * 6 = 9x^2 - 3x * 6 = 9x^2 - 18x

OpenStudy (anonymous):

You got it!

OpenStudy (quickstudent):

:D:D:D f(x) = 56x^8 g(x) = 7x^3 f(x)/g(x) = ? 56x^8/7x^3 = 8x^5

OpenStudy (anonymous):

Correct again.

OpenStudy (quickstudent):

f(x) = 3x g(x) = 3x - 6 f(2) * g(6) = ? (3 * 2) * (3 * 6 - 6) = 6 * (3 * 6 - 6) = 6 * (18 - 6) = 6 * 12 = 72

OpenStudy (anonymous):

Exactly

OpenStudy (quickstudent):

Wow! Thanks! One more left. f(x) = 56x^8 g(x) = 7x^3 f(2)/g(2) = ? (56 * 2^8)/(7 * 2^3) = (56 * 256)/(7 * 2^3) = 14,336/(7 * 2^3) = 14,336/(7 * 8) = 14,336/56 = 256

OpenStudy (anonymous):

Well done!

OpenStudy (quickstudent):

Wow! :D:D:D Thanks for all your help!!! :D